# Large Language Models Can Learn Temporal Reasoning

### 12 Jan 2024 v1

@@ Abstract
Large language models (LLMs) learn temporal concepts from the co-occurrence of related tokens in a sequence. Compared with conventional text generation, temporal reasoning, which reaches a conclusion based on mathematical, logical and commonsense knowledge, is more challenging. In this paper, we propose TempGraph-LLM, a new paradigm towards text-based temporal reasoning. To be specific, we first teach LLMs to translate the context into a temporal graph. A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for pre-training on this task. We prove in experiments that LLMs benefit from the pre-training on other tasks. On top of that, we guide LLMs to perform symbolic reasoning with the strategies of Chain of Thoughts (CoTs) bootstrapping and special data augmentation. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free text.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform temporal reasoning very well as we expected  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). Compared with other singular types, temporal reasoning is a form of integrated task built upon foundations like mathematics, logical and commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
@@
Figure 1: Our TempGraph-LLM framework has two steps: text-to-tempgraph translation and neuro-symbolic reasoning.
@@
Recent work mainly adopts the general methods to investigate and improve the temporal reasoning capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different temporal reasoning tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include span extraction pre-training, instruction tuning and reinforcement learning.
@@
Despite the effectiveness of such methods, they ignore the intrinsic nature of temporal reasoning, and cannot guarantee the correctness of intermediate steps  Saparov and He (2023); Yang et al. (2023b). Human performs temporal reasoning on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline from the text, and then to perform deliberate reasoning on it. However, the high-quality extraction of aligned timeline is a challenging task itself due to lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks.
@@
Given the reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on that. (1) How to introduce the necessary arithmetic and commonsense knowledge involved in temporal reasoning? (2) How to guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap of the timestamps, the relative order of the time gaps). As for (2), there are always two roadmaps: (i) translating natural language into logical statements and using symbolic inference engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design two types of data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
@@
  • We propose a new paradigm, TempGraph-LLM, for text-based temporal reasoning. In this framework, we first translate the context into a temporal graph, and then perform symbolic reasoning on it. Extensive experiments prove that it brings better performance than convetional methods.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLMs to utilize symbolic reasoning, which brings better performance than free text.
  • We present a pipeline to create a synthetic dataet for question answering that requires temporal reasoning. It is fully controllable and requires minimal supervision for context-tempgraph alignment. We show in experiments that pre-training on our dataset benefits LLMs on other tasks.
2 Dataset Construction
We propose a pipeline for the construction of our synthetic temporal QA dataset called TempGraph QA. Each sample, as shown in Table 1, is in the form of (story, tempgraph, question, answer). The tempgraph is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018). The number of events in each tempgraph is restricted into a certain range. By using some rule-based templates, we obtain reliable question and answer pairs from the graph. Finally, each story is translated from the tempgraph by GPT-3.5. Below we show the details for all steps.
TempGraph (sub; rel; obj; start/end; time):
@@ 2 Dataset Construction
Graph-based QA (from Rule-based Python Script):
@@
A: (John Thompson owned Pearl Network).
Q: True or false: event (John Thompson owned Pearl
@@
Table 1: Each sample from TempGraph QA is in the form of (story, tempgraph, question, answer).
@@
Sequencing Which event occurred first, {event A} or {event B}? {event A}/{event B}
Given the following five events: {event A}, {event B}, {event C}, {event D}, {event A}/{event B}/
{event E}, which event is the first/second/third/fourth/fifth one in chronolog- {event C}/{event D}/
ical order? {event E}
Duration How long did the {event A} last? {duration of event A}
Temporal How much time passed between the start of {event A} and the {gap between event A
Relation start of {event B}? and B’s start time}
What happened right before/after {event A} started? {event B} right before/
after {event A}
Facts When did the {event A} occur? {event A’s start time}
@@
Simultaneity True or false: {event A} and {event B} happened at the same year? True / False
True or false: {event A} was still happening when {event B} started? True / False
Comparative True or false: {event A} was longer in duration than {event B}? True / False
@@
Table 2: Reasoning types with corresponding questions and answers in TempGraph QA.
@@
Step 1: Graph Splitting & Masking   Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. To avoid the discrepancy caused by the number of events, we set a certain range and discard the non-compliant subgraphs. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training and test samples. We finally obtain 500 tempgraphs for training and another 100 for test. In addition, we consider the noticeable leakage problem of LLMs, where the prior knowledge of test data has been implicitly obtained from the extensive pre-training. Thus, a masking strategy, i.e., changing entities names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5. To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
@@
Step 2: Graph-based Open QA Creation   Based on the given tempgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones. To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design nine different question types (shown in Table 2), and write a rule-based Python script to determine the answers from the tempgraph.
@@
Step 3: Quality Control   Although LLMs are good at story generation, there is a risk of misalignment between the given tempgraph and the generation text. To avoid the interference caused by misalignment, we propose an automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we remove the possibly missing events in the graph. Since this strategy tends to simplify the reasoning, we manually verify the existence of the possibly missing events for both training and test sets. Note that only the quality of story-tempgraph alignment needs control, since all the QAs are generated from the tempgraph with specific rules. We show dataset statistcis in Appendix A, and all the prompts involved in Appendix C.
3 TempGraph-LLM
@@
Motivated by human perception, we propose a new paradigm called TempGraph-LLM. We first translate the text into a temporal graph, and then guide LLMs to perform symbolic reasoning on it.
3.1 Text-to-TempGraph Translation
@@
TempGraph (TG) serves as the foundation of the following symbolic reasoning process. Although LLMs can extract them from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs. Thus, we provide a deliberate pipeline for high-quality TG translation.
3.1.1 Pre-training
In the TempGraph QA, each story is corresponding to a verified TG. Similarly, for other data, where the ground truth of TG is avaiable, we can pre-train LLMs on them. Since language models (LMs) can only deal with sequence of tokens, we sort all the events in the graph according to their time. In addition, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the pre-training on our data for other tasks.
3.1.2 Fine-tuning
@@
Generally, the biggest challenge for fine-tuning on real-world datasets is lack of ground truth. Instead of directly using the estimation from LLMs, we provide a pipeline for better alignment.
@@
Entity & Relation Extraction   To align the entities and relations in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entities and relations in pre-defined QAs, we consider two ways: parsing the QAs with rules or using LLMs. The former is more efficient and reliable when the QAs are generated by templates  (Chen et al., 2021; Tan et al., 2023a), while the latter should be used when the QAs are in free text  (Zhou et al., 2019; Ning et al., 2020).
@@
Temporal Info Identification   Compared with conventional reasoning tasks, temporal information is one of the most important and unique components here. To identify all the time expressions, LLMs are first used to extract them in the context, and then a rule-based Python script  (Yang et al., 2023a) is adopted for time-normalization and invalid output filtering. For example, we convert "Dec. 5th, 2004" into "2004-12-05", "July 1989" into "1989-07-XX", and "1973" into "1973-XX-XX", respectively. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, e.g., "1974 is before 1983", "the time gap between May 2004 and Oct 2008 is 4 years and 5 months", which are introduced in temporal reasoning as external knowledge.
@@
Graph Construction & Verification   We generate the graph using LLMs with in-context learning (ICL). To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output. The TG should only contain the extracted elements to guarantee the alignment with pre-defined QAs. We classify the stories according to whether they contain explicit expressions such as "July 1989", "2003–2005", "the late 1900s". For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
@@ 3.2 Neuro-symbolic Temporal Reasoning
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by Chain of Thoughts (CoTs) bootstrapping and data augmentation.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples. We bootstrap the generations and only accept the ones with correct final answers.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{m}\}$ with LLMs. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask LLMs to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c^{k}_{j^{\prime}},\hat{a}^{k}_{j^{\prime}})$ from LLMs, and only accept the one leading to correct answers, i.e., $\hat{a}^{k}_{j^{\prime}}=a^{*}_{j^{\prime}}$. Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method is more flexible and robust. In addition, inspired by  (Wang et al., 2023), we compare different type of CoTs using contrastive learning. The core idea is to choose CoTs that make the correct answer more plausible than other candidates.
$c^{*}=\operatorname{argmax}\log P\left(a^{*}\mid g,e,q,c\right)+\gamma G\left(c\right)$ (1)
$G\left(c\right)=\log\frac{P\left(a^{*}\mid g,e,q,c\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c\right)}$ (2)
$P\left(a\mid g,e,q,c\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c,t_{<l}\right)$ (3)
where $g,e,q,c$ denote the TG, external knowledge, question, and CoTs, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.

### 20 Feb 2024 v2

@@ Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate contextual details. In this paper, we propose TG-LLM, a new framework towards language-based TR. To be specific, we first teach LLM to translate the context into a temporal graph (TG). A synthetic dataset, which is fully controllable and requires minimal supervision, is constructed for fine-tuning on this graph translation task. We confirm in experiments that the capability of TG extraction learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we guide LLM to perform symbolic reasoning over the TG via Chain of Thoughts (CoTs) bootstrapping and special data augmentation strategies. We observe that CoTs with symbolic reasoning bring more consistent and reliable results than those using free-form text.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as order or duration; it extends to more intricate aspects, e.g., making sense of complex narratives or discovering causal relations. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have shown remarkable performance in complex reasoning tasks  (Huang and Chang, 2023). However, they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023). TR is more challenging since it requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
@@
Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and neuro-symbolic reasoning.
@@
Recent works mainly adopt the general reasoning methods to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark leading LLMs on different TR tasks with standard prompting, few-shot learning and Chain of Thoughts (CoTs)  (Wei et al., 2023a).  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). Similarly,  (Wei et al., 2023b) design several special types of prompts for in-context learning (ICL). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Despite the effectiveness of such methods, they either ignore or not explicitly involve the intrinsic nature of TR. Human performs complex TR on a timeline of events which are aligned with the entities and relations. These temporal concepts (e.g., ordering, duration, frequency, typical time) are then rigorously defined based on the timeline information. Thus, a natural solution for LLMs to reach human-level performance is to first extract a timeline of events from the text, and then perform deliberate reasoning on it. However, the high-quality extraction of the aligned timeline is a challenging task itself due to the lack of ground truth. To solve this problem, we propose a synthetic dataset, which is fully controllable and requires minimal supervision. We show in experiments that the capability of timeline extraction learned on our dataset can be transferred to other tasks and benchmarks.
@@
Given a reliable timeline, we still need to solve two problems to guide LLMs to perform rigorous reasoning on it: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? (2) How can one guide LLMs to perform rigorous reasoning? Inspired by retrieval-augmented generation (RAG)  (Lewis et al., 2021), we first extract all the time expressions, and then generate all the necessary information (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps). As for (2), there are two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) directly using LLMs as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of the generated text especially when training data are insufficient  (Yang et al., 2023b). To solve these problems, we propose a symbolic-integration bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. Furthermore, we design special data augmentation strategies to mitigate data deficiency for neuro-symbolic reasoning tasks.
@@
  • We propose a new paradigm, TG-LLM, for language-based TR. In this framework, we first translate the context into a temporal knowledge graph, and then perform symbolic reasoning on it. Extensive experiments prove that our novel approach results in superior performance compared to the conventional approach.
  • We design two approaches including chain of thoughts bootstrapping and data augmentation to guide LLM to utilize symbolic reasoning, which brings better performance than free-form text.
  • We present a pipeline to create a synthetic dataset for question answering that requires TR. It is fully controllable and requires minimal supervision for text-temporal graph alignment. We show in experiments that fine-tuning on our dataset benefits LLMs on other TR tasks and benchmarks.
Temporal Graph (sub; rel; obj; start/end; time):
@@
Graph-based QAs (from rule-based Python script):
@@
A1: (John Thompson owned Pearl Network).
Q2: True or false: event (John Thompson owned Pearl
@@
Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
2 Dataset Construction
@@
Sequencing Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_A>/<Event_B>/
<Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in <Event_C>/<Event_D>/
the chronological order? $\cdots$
Duration How long did the event <Event_A> last? <Duration_of_Event_A>
Temporal How much time passed between the start of <Event_A> and the start of <Gap_between_Event_A
Relation <Event_B>? _and_B_start_time>
What happened right before/after <Event_A> started? <Event_B> right before/
after <Event_A>
Facts When did the <Event_A> occur? <Event_A_start_time>
@@
Simultaneity True or false: <Event_A> and <Event_B> happened at the same year? True / False
True or false: <Event_A> was still happening when <Event_B> started? True / False
Comparative True or false: <Event_A> was longer in duration than <Event_B>? True / False
@@
Table 2: Reasoning types with corresponding questions and answers in TGQA.
We propose a pipeline for the construction of our synthetic dataset called TGQA. Each sample, as shown in Table 1, is in the form of (story, temporal graph, questions, answers). The temporal graph (TG) is a subgraph extracted from a large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), with a restriction on the number of events. Each story is translated from the TG by GPT-3.5. Finally, by using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. We obtain 500 samples for training/validation and another 100 for test, with about 30 QA pairs in a single sample. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of temporal questions as shown in Table 2. Below we provide the details of each step in the dataset construction pipeline.
@@
Step 1: Graph Splitting & Masking  Existing temporal knowledge graph datasets (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) have tens of thousands of events, making it impossible to convert them into text all at once. Given a specific entity, we extract its third-order neighbors and all the events happening between them. Since we want LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the facts in training/validation and test samples. In addition, we consider the noticeable corpus leakage problem of LLMs, where prior knowledge of test data has been implicitly obtained from the extensive pre-training process. Thus, a masking strategy, i.e., changing entity names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5 (see Table 6). To avoid confusion, we only choose obscure names that do not exist in the original large knowledge graph.
@@
Step 2: Graph-based Open QA Creation  Based on the given TG, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the graph with the requirements to include all the facts. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we split the starting and ending of a persistent event into two transient ones (see Table 1). To obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple diverse question types (see Table 2), and use a rule-based Python script to generate the corresponding Qs and As.
@@
Step 3: Quality Control  Although LLMs are good at story generation, there is a risk of misalignment between the given TG and the generated text. To avoid the interference caused by misalignment, we propose a semi-automatic way for verification. Given a generated story, we ask LLMs the time of each event in the graph. If LLMs cannot give the correct answer, we consider the event as possibly missing in the story. Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors and only manually inspect the set of unanswered questions by LLM. Note that only the quality of story-TG alignment needs manual inspection, since all the QAs are generated from the TG with specific rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
3 TG-LLM
@@
Motivated by human perception, we propose a new paradigm called TG-LLM. We first translate the text into a temporal graph, and then guide the LLM to perform symbolic reasoning on it.
3.1 Text-to-TG Translation
@@
TG serves as the foundation of the following symbolic reasoning process. Although LLMs can extract fact/events from the given context without fine-tuning, we observe a misalignment between the generated graph and pre-defined QAs, i.e, LLMs making mistakes or using different entity/relation names or time expressions. Thus, we provide a deliberate pipeline for high-quality TG translation.
@@
Ground-truth TG Generation   For datasets such as TGQA, each story is corresponding to a verified TG. Thus, we can directly fine-tune LLMs on this task with the ground truth. However, the biggest challenge for most real-world applications is the lack of ground truth. Instead of using the naive estimation from LLMs, we provide a pipeline for high-quality generation.
@@
(1) Entity & Relation Extraction: To align the entity/relation names in generated TGs and pre-defined QAs, we use a prompt with requirements on including the ones in pre-defined QAs when extracting from the context. To obtain the entity/relation names in pre-defined QAs, we consider two ways: parsing with rules or using LLMs.
@@
(2) Temporal Info Identification: To identify all the time expressions, LLMs are first used to extract them from the context, and then a rule-based Python script is adopted for time normalization and invalid output filtering. For the valid normalized time expressions, we can easily obtain their temporal relations and time gaps, which are introduced as external knowledge in temporal reasoning (see Table 7).
@@
(3) Graph Construction & Verification: We generate the graph using LLMs with ICL. To be specific, we provide in the prompt several examples which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (see Table 15). For those without explicit expressions, we ask LLMs to provide the temporal order of different events. We further semi-automatically verify the generated TGs by asking LLMs to answer the pre-defined QAs based on the graph. For those graphs where LLMs give the wrong answers, we manually check and correct them.
Figure 2: In chain of thoughts (CoTs) bootstrapping, we accept CoTs leading to correct answers and sample them according to their contrastive learning scores.
Supervised Fine-tuning  We found out in our experiments that providing a chronological list of events helps LLMs provide better temporal reasoning. In addition, as shown in Table 1, to deal with both transient and persistent events, we consider the starting and ending of a persistent event as two transient events. With these steps, we transfer the TG into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
@@ 3.2 Neuro-symbolic Temporal Reasoning
Given the reliable TGs, we train LLMs to perform symbolic reasoning with SFT enhanced by CoTs bootstrapping and data augmentation.
@@ 3.2.1 Bootstrapping Chain of Thoughts
It is observed that SFTs on reliable CoTs bring better performance in reasoning tasks than those on standard Input/Output prompts  (Wang et al., 2023; Ho et al., 2023). However, it is challenging to obtain reliable CoTs from LLMs for symbolic reasoning tasks  (Saparov and He, 2023). Inspired by  (Zhu et al., 2023), we provide several manually verified examples to LLMs as ICL, and ask LLMs to generate CoTs for the new samples.
To be specific, we first classify the pre-defined QAs into several categories $\{Q_{1},Q_{2},\cdots,Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$ and ask GPT-3.5/GPT-4 to provide symbolic CoTs $\{c_{j}\}$ given the TG $\{g_{j}\}$, external knowledge $\{e_{j}\}$ and question $\{q_{j}\}$. These CoTs will be manually verified and used as few-shot ICL examples. When given new samples $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}},a_{j^{\prime}})$, we bootstrap CoTs with final answers $(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})$ from GPT-3.5, and only accept the one leading to correct answers, i.e., $\hat{a}_{j^{\prime},k}=a^{*}_{j^{\prime}}$. In addition, inspired by  (Wang et al., 2023), we compare different CoTs using contrastive learning given in Eq. 2. The core idea is to choose CoTs that make the correct answer more plausible than other candidate answers. For the generated CoTs, we use weighted sampling according to the score-based probability given in Eq. 1.
$P_{sample}(c_{k})=\operatorname{softmax}(score(c_{k}))$ (1)
$score(c_{k})=\log P\left(a^{*}\mid g,e,q,c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid g,e,q,c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c_{k}\right)}$ (3)
$P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
where $P_{sample}(\cdot)$ is the sampling probability for CoTs selection, $score(\cdot)$ is the contrastive learning score for each CoT, $g,e,q,c$ denote the TG, external knowledge, question, and CoT, respectively; $a^{*}$ denotes the correct answer; $a^{\prime}$ denotes the wrong answer from a candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $G(\cdot)$ denotes the contrastive plausibility growth; conditional probability $P(\cdot)$ is calculated in a way similar to perplexity, $t_{l}$ denotes the $l$-th token in the answer $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
Compared with generating CoTs with templates  (Saparov and He, 2023; Yuan et al., 2023), our method has more diversity and does not require prior knowledge for the reasoning process.